In an LR-circuit, the inductive reactance is equal to the resistance R of the circuit. An emf E=E0cos(wt) applied to the circuit. The power consumed in the circuit is
A) E02/R
B)E02/2R
C)E02/4R
D)E02/8R

Grade:

2 Answers

Sanjai
4 years ago
 XL= R.       Z= 2 R
   P= ( Vrms/ Z)•R
     = ( Eo/ √2)2 / (√2R)2 • R
     = Eo 2/4R
 
please adjust with the symbols. I couldn’t write it clearly.
 
      
     
Khimraj
4 years ago
  •  P={{E}_{rms}}{{i}_{rms}}\cos \varphi =\frac{{{E}_{0}}}{\sqrt{2}}\times \frac{{{i}_{0}}}{\sqrt{2}}\times \frac{R}{Z}                    Þ \frac{{{E}_{0}}}{\sqrt{2}}\times \frac{{{E}_{0}}}{Z\sqrt{2}}\times \frac{R}{Z}\Rightarrow \,\,P=\frac{E_{0}^{2}R}{2{{Z}^{2}}}            Given {{X}_{L}}=R so, Z=\sqrt{2}R\Rightarrow \,P=\frac{E_{0}^{2}}{4R}